An object of mass 0.01 kg is moving in a circular path of radius 150 cm. if it takes 3 sec to make one complete revolution. calculate the magnitude and the direction of the centripetal force.


1
Expert's answer
2020-06-12T11:16:11-0400

The centripetal force will be


"F=m\frac{v^2}{R}=m\frac{(2\pi R/T)^2}{R}=\frac{4m\pi^2R}{T},\\\space\\
F=\frac{4\cdot0.01\pi^2\cdot1.5}{3}=0.20\text{ N}."

The force is directed into the center of the circle perpendicularly to the object's velocity.